# Phase transitions for unique codings of fat Sierpinski gasket with multiple digits

### v1

@@ 4.2. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
By the definition of $p_{i}$ in (4.18) it follows that $p=p_{4}\geq 5$. So, by (4.17) Propositions 2.3 and 4.2 it follows that $d_{m+p-4}=\alpha_{N(N+1)}$ and $d_{m+p-5}=\alpha_{(N+1)N}$ or $\alpha_{NN}$. However, if $d_{m+p-5}=\alpha_{(N+1)N}$, then by Proposition 4.2 we have
@@
which implies $p=p_{4}\geq 6$ by (4.18). By (4.17), Propositions 2.3 and 4.2 we obtain that $d_{m+p-5}d_{m+p-4}=\alpha_{N(N+1)}\alpha_{(N+1)N}$ and $d_{m+p-6}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{(N+1)(N-1)}\right\}$. Similar to Claim 2 we can prove that $d_{m+p-6}\neq\alpha_{(N+1)(N-1)}$.
@@
for all $1\leq i\leq\lfloor\frac{p}{3}\rfloor$. This leads to a contradiction with (4.18), completing the proof. ∎
@@
Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. If $M=3N+1$ with $N=0$, then it is clear that $d_{i}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{N(N+1)}\right\}$ for all $i\geq 1$. In the following we assume $N\geq 1$. Note by Proposition 3.1 that $\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\widetilde{{\mathbf{U}}}_{\beta_{G},M}=\left\{\alpha_{0M}^{\infty},\alpha_{M0}^{\infty},\alpha_{00}^{\infty}\right\}$. Furthermore, by Proposition 4.2 and Lemma 2.1 that $\delta(\beta_{c})\prec(N+1)^{\infty}$. Then there exists $m_{0}\in\mathbb{N}$ such that $d_{m_{0}}^{1}<M,d_{m_{0}}^{1}<M$ and $\overline{d_{m_{0}}^{\oplus}}<M$. Since $M=3N+1$ with $N\geq 1$, by (2.1) and Proposition 2.3 it follows that
@@
  • Note by the definition of $(\mathbf{a}_{n})$ in (4.14). Then Proposition 4.7 implies that if $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\mathbf{a}_{n}^{+}$, then $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}$ or $\mathbf{a}_{n+1}^{+}$. Similarly, if $d_{m}^{1}>0$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\Phi_{1}(\mathbf{a}_{n}^{+})$, then $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\Phi_{1}(\mathbf{a}_{n+1})$ or $\Phi_{1}(\mathbf{a}_{n+1}^{+})$.
@@
However, in the first case we have $d_{7}^{2}=N<M$ and $d_{8}^{2}d_{9}^{2}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$; and in the second case we have $\overline{d_{7}^{\oplus}}=N<M$ and $\overline{d_{8}^{\oplus}d_{9}^{\oplus}}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$. Both cases lead to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+4}=\alpha_{NN}$.
@@
This gives $\overline{d_{m+7}^{\oplus}}=N<M$ and $\overline{d_{m+8}^{\oplus}d_{m+9}^{\oplus}}=(N+1)(N+1)$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+7}=\alpha_{NN}$.
@@
Since the proof of (ii) is similar, we only prove (i). This will be done by induction on $n$. Note by Lemmas 4.9 and 4.10 that (i) holds for $n=0,1$. Now we assume (i) holds for all $k\leq n$ for some $n\in\mathbb{N}$, and we will prove it for $k=n+1$. Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. As in the proof of Lemma 4.9 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 4.6 there exists $m_{1}\in\mathbb{N}$ such that $d_{j}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $j>m_{1}$. Now take $m>m_{1}$, and suppose $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}^{+}$ with $d_{m}^{1}<M$. We will prove
@@
Observe by (4.14) that $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}^{+}=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n})=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n-1}^{+})\mathbf{a}_{n-1}^{+},$ which gives that $d^{1}_{m+3(2^{n}+2^{n-1})}=N<M$ and $d_{m+3(2^{n}+2^{n-1})+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n-1}^{+}$. Then by the induction hypothesis for $k=n-1$ we have
@@ 4.3. Phase transition for $\widetilde{{\mathbf{U}}}_{\beta,M}$ with $M=3N+1$
Recall that the sequence $(\beta_{n})$ is defined in (4.12). Note that $\beta_{n}\nearrow\beta_{c}$ as $n\to\infty$. Our strategy on proving Theorem 1.4 (ii) is to show that the difference set $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ is countably for any $n\geq 0$.
@@
(4.25) $d_{m+1}d_{m+2}=\alpha_{(N+1)N}\alpha_{N(N+1)}\quad\textrm{and}\quad d_{m+3}\in\left\{\alpha_{(N+1)N},\alpha_{NN}\right\};$
@@
leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. So, $d_{m+4}d_{m+5}d_{m+6}=\Phi_{1}(\mathbf{a}_{0}^{+})$. Again, by Proposition 4.7 and the same argument as above we can deduce that
@@
However, by (4.15) this implies that $d^{1}_{m+1}d^{1}_{m+2}\ldots={\mathbf{t}}_{1}^{\infty}=\delta(\beta_{1})$, which again leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$.
@@
We will prove the theorem by induction on $n$. Note by Lemma 4.12 that it holds for $n=0$. Now suppose the result holds for all $n\leq k$ for some $k\geq 0$, and we consider it for $n=k+1$. Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$. Then by Lemma 4.6 there exists $m_{1}\in\mathbb{N}$ such that $d_{m_{1}}^{1}<M$ and $d_{m_{1}+i}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $i\geq 1.$ Note by the induction hypothesis that any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\bigcup_{j=0}^{k}(\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M})$ must end with $\mathbf{a}_{j}^{\infty},\Phi_{1}(\mathbf{a}_{j}^{\infty})$ or their cyclic permutations under $\theta$ for some $0\leq j\leq k$. Observe by (4.15) that $\mathbf{a}_{j}^{1}={\mathbf{t}}_{j}$ for all $j\geq 0$, and ${\mathbf{t}}_{0}^{\infty}\prec{\mathbf{t}}_{1}^{\infty}\prec\cdots\prec{\mathbf{t}}_{k}^{\infty}\prec{\mathbf{t}}_{k+1}^{\infty}$. Since $\delta(\beta_{j})={\mathbf{t}}_{j}^{\infty}$ for all $j\geq 0$, by (4.14) it follows that the sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$ must contain a block in $\bigcup_{j=0}^{2}\left\{\theta^{j}(\mathbf{a}_{k}^{+}),\theta^{j}(\Phi_{1}(\mathbf{a}_{k}^{+}))\right\}$. Without loss of generality we may assume $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k}}=\mathbf{a}_{k}^{+}$ for some $m_{2}\geq m_{1}$. Note that $d^{1}_{m_{2}}<M$. Then by Proposition 4.7 it follows that
@@
If $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k+1}}=\mathbf{a}_{k+1}^{+}$, the by Proposition 4.7 and using $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$ it follows that $d_{m_{2}+1}d_{m_{2}+2}\ldots=\mathbf{a}_{k+2}^{\infty}$. But by (4.15) this gives
@@
which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
@@
If $d_{m_{2}+3\cdot 2^{k+1}+1}\ldots d_{m_{2}+3(2^{k+1}+2^{k})}=\mathbf{a}_{k}$, then $d_{m_{2}+3\cdot 2^{k}+1}\ldots d_{m_{2}+3\cdot(2^{k+1}+2^{k})}=\Phi_{1}(\mathbf{a}_{k}^{+})\mathbf{a}_{k}=\Phi(\mathbf{a}_{k+1}^{+})$. By (4.15), Proposition 2.3 and using $d^{1}_{m_{2}+3\cdot 2^{k}}>0$ it follows that
@@
again leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
@@
First we prove (ii). Take $\beta\in(\beta_{G},\beta_{c})$. Note that $\beta_{0}=\beta_{G}$ and $\beta_{k}\nearrow\beta_{c}$ as $k\to\infty$. So there exists $k\in\mathbb{N}$ such that $\beta<\beta_{k}$. Note by Propositions 3.1 and 4.11 that $\widetilde{{\mathbf{U}}}_{\beta_{k},M}=\widetilde{{\mathbf{U}}}_{\beta_{0},M}\cup\bigcup_{j=0}^{k-1}\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M}$ is at most countable. This implies that $\widetilde{{\mathbf{U}}}_{\beta,M}\subset\widetilde{{\mathbf{U}}}_{\beta_{k},M}$ is at most countable. On the other hand, by Proposition 2.3 we have
@@
Next we prove (iii). The proof is similar to [23, Proposition 5.6]. Note that $\beta_{k}\nearrow\beta_{c}$ as $k\to\infty$. By Proposition 4.7 it follows that $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ contains all of the following sequences
@@
=$\mathscr{F}=\left\{\alpha_{10}\alpha_{10},\alpha_{01}\alpha_{01},\alpha_{00}\alpha_{00}\right\}.$
@@
We remark that if $\beta$ is a mutinacci number, then by a similar argument as in Example 4.13 we are able to determine the explicit Hausdorff dimension of $U_{\beta,1}$.

### v2

@@ 4.2. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
By the definition of $p_{i}$ in (4.18) it follows that $p=p_{4}\geq 5$. So, by (4.17), and Propositions 2.3 and 4.2 it follows that $d_{m+p-4}=\alpha_{N(N+1)}$ and $d_{m+p-5}=\alpha_{(N+1)N}$ or $\alpha_{NN}$. However, if $d_{m+p-5}=\alpha_{(N+1)N}$, then by Proposition 4.2 we have
@@
which implies $p=p_{4}\geq 6$ by (4.18). By (4.17), and Propositions 2.3 and 4.2 we obtain that $d_{m+p-5}d_{m+p-4}=\alpha_{N(N+1)}\alpha_{(N+1)N}$ and $d_{m+p-6}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{(N+1)(N-1)}\right\}$. Similar to Claim 2 we can prove that $d_{m+p-6}\neq\alpha_{(N+1)(N-1)}$.
@@
for all $1\leq i\leq\lfloor\frac{p}{3}\rfloor$. This leads to a contradiction to (4.18), completing the proof. ∎
@@
Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. If $M=3N+1$ with $N=0$, then it is clear that $d_{i}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{N(N+1)}\right\}$ for all $i\geq 1$. In the following we assume $N\geq 1$. Note by Proposition 3.1 that $\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\widetilde{{\mathbf{U}}}_{\beta_{G},M}=\left\{\alpha_{0M}^{\infty},\alpha_{M0}^{\infty},\alpha_{00}^{\infty}\right\}$. Furthermore, by Proposition 4.2 and Lemma 2.1 that $\delta(\beta_{c})\prec(N+1)^{\infty}$. Then there exists $m_{0}\in\mathbb{N}$ such that $d_{m_{0}}^{1}<M,d_{m_{0}}^{2}<M$ and $\overline{d_{m_{0}}^{\oplus}}<M$. Since $M=3N+1$ with $N\geq 1$, by (2.1) and Proposition 2.3 it follows that
@@
  • Recall $(\mathbf{a}_{n})$ from (4.14). Then Propositin 4.8 implies that if $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\mathbf{a}_{n}^{+}$, then $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}$ or $\mathbf{a}_{n+1}^{+}$. Similarly, if $d_{m}^{1}>0$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\Phi_{1}(\mathbf{a}_{n}^{+})$, then $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\Phi_{1}(\mathbf{a}_{n+1})$ or $\Phi_{1}(\mathbf{a}_{n+1}^{+})$.
@@
However, in the first case we have $d_{7}^{2}=N<M$ and $d_{8}^{2}d_{9}^{2}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$; and in the second case we have $\overline{d_{7}^{\oplus}}=N<M$ and $\overline{d_{8}^{\oplus}d_{9}^{\oplus}}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$. Both cases lead to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+4}=\alpha_{NN}$.
@@
This gives $\overline{d_{m+7}^{\oplus}}=N<M$ and $\overline{d_{m+8}^{\oplus}d_{m+9}^{\oplus}}=(N+1)(N+1)$, leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+7}=\alpha_{NN}$.
@@
Since the proof of (ii) is similar, we only prove (i). This will be done by induction on $n$. Note by Lemmas 4.10 and 4.11 that (i) holds for $n=0,1$. Now we assume (i) holds for all $k\leq n$ with $n\in\mathbb{N}$, and we will prove it for $k=n+1$. Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. As in the proof of Lemma 4.10 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 4.7 there exists $m_{1}\in\mathbb{N}$ such that $d_{j}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $j>m_{1}$. Now take $m>m_{1}$, and suppose $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}^{+}$ with $d_{m}^{1}<M$. We will prove
@@
Observe by (4.14) that $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}^{+}=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n})=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n-1}^{+})\mathbf{a}_{n-1}^{+},$ which gives that $d^{1}_{m+3(2^{n}+2^{n-1})}=N<M$ and $d_{m+3(2^{n}+2^{n-1})+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n-1}^{+}$. Then by the induction hypothesis for k=n1k=n-1 we have
@@ 4.3. Phase transition for $\widetilde{{\mathbf{U}}}_{\beta,M}$ with $M=3N+1$
Recall that the sequence $(\beta_{n})$ is defined in (4.12). Note that $\beta_{n}\nearrow\beta_{c}$ as $n\to\infty$. Our strategy on proving Theorem 1.4 (ii) is to show that the difference set $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ is countable for any $n\geq 0$.
@@
(4.25) $d_{m+1}d_{m+2}=\alpha_{(N+1)N}\alpha_{N(N+1)}\quad\textrm{and}\quad d_{m+3}\in\left\{\alpha_{(N+1)N},\alpha_{NN}\right\},$
@@
leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. So, $d_{m+4}d_{m+5}d_{m+6}=\Phi_{1}(\mathbf{a}_{0}^{+})$. Again, by Proposition 4.8 and the same argument as above we can deduce that
@@
However, by (4.15) this implies that $d^{1}_{m+1}d^{1}_{m+2}\ldots={\mathbf{t}}_{1}^{\infty}=\delta(\beta_{1})$, which again leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$.
@@
We will prove the theorem by induction on $n$. Note by Lemma 4.13 that it holds for $n=0$. Now suppose the result holds for all $n\leq k$ with $k\geq 0$, and we consider it for $n=k+1$. Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$. Then by Lemma 4.7 there exists $m_{1}\in\mathbb{N}$ such that $d_{m_{1}}^{1}<M$ and $d_{m_{1}+i}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $i\geq 1.$ Note by the induction hypothesis that any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\bigcup_{j=0}^{k}(\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M})$ must end with $\mathbf{a}_{j}^{\infty},\Phi_{1}(\mathbf{a}_{j}^{\infty})$ or their cyclic permutations under $\theta$ for some $0\leq j\leq k$. Observe by (4.15) that $\mathbf{a}_{j}^{1}={\mathbf{t}}_{j}$ for all $j\geq 0$, and ${\mathbf{t}}_{0}^{\infty}\prec{\mathbf{t}}_{1}^{\infty}\prec\cdots\prec{\mathbf{t}}_{k}^{\infty}\prec{\mathbf{t}}_{k+1}^{\infty}$. Since $\delta(\beta_{j})={\mathbf{t}}_{j}^{\infty}$ for all $j\geq 0$, by (4.14) it follows that the sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$ must contain a block in $\bigcup_{j=0}^{2}\left\{\theta^{j}(\mathbf{a}_{k}^{+}),\theta^{j}(\Phi_{1}(\mathbf{a}_{k}^{+}))\right\}$. Without loss of generality we may assume $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k}}=\mathbf{a}_{k}^{+}$ for some $m_{2}\geq m_{1}$. Note that $d^{1}_{m_{2}}<M$. Then by Proposition 4.8 it follows that
@@
If $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k+1}}=\mathbf{a}_{k+1}^{+}$, then by Proposition 4.8 and using $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$ it follows that $d_{m_{2}+1}d_{m_{2}+2}\ldots=\mathbf{a}_{k+2}^{\infty}$. But by (4.15) this gives
@@
which leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
@@
If $d_{m_{2}+3\cdot 2^{k+1}+1}\ldots d_{m_{2}+3(2^{k+1}+2^{k})}=\mathbf{a}_{k}$, then $d_{m_{2}+3\cdot 2^{k}+1}\ldots d_{m_{2}+3\cdot(2^{k+1}+2^{k})}=\Phi_{1}(\mathbf{a}_{k}^{+})\mathbf{a}_{k}=\Phi(\mathbf{a}_{k+1}^{+})$. By (4.15), and Proposition 2.3 and using $d^{1}_{m_{2}+3\cdot 2^{k}}>0$ it follows that
@@
again leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
@@
First we prove (ii). Take $\beta\in(\beta_{G},\beta_{c})$. Note that $\beta_{0}=\beta_{G}$ and $\beta_{k}\nearrow\beta_{c}$ as $k\to\infty$. So there exists $k\in\mathbb{N}$ such that $\beta<\beta_{k}$. Note by Propositions 3.1 and 4.12 that $\widetilde{{\mathbf{U}}}_{\beta_{k},M}=\widetilde{{\mathbf{U}}}_{\beta_{0},M}\cup\bigcup_{j=0}^{k-1}\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M}$ is at most countable. This implies that $\widetilde{{\mathbf{U}}}_{\beta,M}\subset\widetilde{{\mathbf{U}}}_{\beta_{k},M}$ is at most countable. On the other hand, by Proposition 2.3 we have
@@
Next we prove (iii). The proof is similar to that of [24, Proposition 5.6]. Note that $\beta_{k}\nearrow\beta_{c}$ as $k\to\infty$. By Proposition 4.8 it follows that $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ contains all of the following sequences
@@
$\mathscr{F}=\left\{\alpha_{10}\alpha_{10},\alpha_{01}\alpha_{01},\alpha_{00}\alpha_{00}\right\}.$
@@
We remark that if $\beta$ is a multinacci number, then by a similar argument as in Example 4.14 we are able to determine the explicit Hausdorff dimension of $U_{\beta,1}$.